# Two variants of the friendship paradox: The condition for inequality between them

### v1

@@ Abstract
The friendship paradox—the observation that, on average, one’s friends have more friends than oneself—admits two common formulations depending on whether averaging is performed over edges or over nodes. These two definitions, the “alter-based” and “ego-based” means, are often treated as distinct but related quantities. This document establishes their exact analytical relationship, showing that the difference between them is governed by the degree-degree covariance $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})$ normalized by the mean degree $\langle k\rangle$. Explicit examples demonstrate the three possible cases of positive, zero, and negative covariance, corresponding respectively to assortative, neutral, and disassortative mixing patterns. The derivation further connects the covariance form to the moment-based expression introduced by Kumar, Krackhardt, and Feld [Proc. Natl. Acad. Sci. 121, e2306412121 (2024)], which involves the (1)(-1)st, 11st, 22nd, and 33rd moments of the degree distribution. The two formulations are shown to be equivalent, as they should be: the moment-based representation expands the same structural dependence that the covariance form expresses in its most compact and interpretable form. The analysis thus unifies node-level and moment-level perspectives on the friendship paradox, offering both a pedagogically transparent derivation and a direct bridge to recent theoretical developments.
@@ I Introduction
Recent work by Kumar, Krackhardt, and Feld [2] introduced a moment-based decomposition that expresses the difference between these two averages in terms of degree moments and an edge-level correlation coefficient, the inversity $\rho$, which captures the correlation between origin degree and inverse destination degree across oriented edges. The present note establishes a compact node-level covariance identity that expresses the difference between the alter-based and ego-based averages as a normalized covariance between degree and neighbor-averaged degree. The equivalence between the two formulations is shown explicitly, including a clear mapping between node-averaged quantities, edge-sampled expectations, and degree moments. This connection illuminates the roles of heterogeneity and mixing patterns [3] and provides a concise criterion for equality of the two variants.
@@ VI Discussion and closing remarks
Epilogue
The configuration shown in Fig. 1 was originally conceived as a question for an undergraduate network science midterm, where students were asked to compute either $\langle k_{\text{friend}}\rangle_{\mathrm{n}}$ or $\langle k_{nn}\rangle_{\mathrm{n}}$. During verification, the unexpected equality $\langle k_{\text{friend}}\rangle_{\mathrm{n}}=\langle k_{nn}\rangle_{\mathrm{n}}$ revealed the existence of the neutral case $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})=0$. This serendipitous observation motivated the analytical derivation of the general inequality condition presented here. Although the simplicity of the covariance expression suggests that the result may not warrant an independent research publication in a refereed journal, it provides a pedagogically elegant synthesis linking structural assortativity, moment relations, and the two canonical definitions of the FP.
The equivalence relation was first identified with the assistance of ChatGPT 5, which also contributed to the final algebraic derivation of the covariance formulation. The author, a non-native English speaker, expresses sincere appreciation to ChatGPT 5 for its help in improving the readability of the manuscript, allowing the ideas to be presented with greater clarity and stylistic polish in a short time. On a cautionary note, though, ChatGPT 5 also made a number of sloppy mistakes in the mathematical derivation throughout the process, so the author has carefully checked every detail and completed the manuscript.
Remark. This note may also be viewed as an experiment in scientific communication. It explores whether concise, rapidly produced expository works—developed with substantial assistance from large language models (LLMs)—can serve as an efficient and legitimate complement to the traditional journal publication process. Although there remain many practical and institutional reasons for pursuing formal peer-reviewed publication, the possibility of disseminating short analytical results or conceptual clarifications directly through platforms such as arXiv could enhance the accessibility and timeliness of research communication. Whether or not this particular note passes arXiv’s screening policies on LLM-assisted writing, the author hopes it can serve as a modest example of how human-AI collaboration might facilitate the clear and rapid sharing of scientific insights, and perhaps as a small guidepost for students or researchers exploring related aspects of network science.
@@ Acknowledgements.
The author sincerely appreciates Hang-Hyun Jo (조항현) for verifying the covariance formula and for his encouragement to share this note publicly on arXiv. The author also thanks Eun Lee (이은), with whom he initially discussed the problem and explored a plan to extend the analysis, as well as Mi Jin Lee (이미진) for organizing an informal workshop dedicated to the discussion of “incomplete ideas” where the topic was further developed. This research was supported by the National Research Foundation of Korea (NRF) under the grant RS-2021-NR061247.

### v2

@@ Abstract
The friendship paradox—the observation that, on average, one’s friends have more friends than oneself—admits two common formulations depending on whether averaging is performed over edges or over nodes. These two definitions, the “alter-based” and “ego-based” means, are often treated as distinct but related quantities. This paper establishes their exact analytical relationship, showing that the difference between them is governed by the degree-degree covariance $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})$ normalized by the mean degree $\langle k\rangle$. Explicit examples demonstrate the three possible cases of positive, zero, and negative covariance, corresponding respectively to assortative, neutral, and disassortative mixing patterns. The derivation further connects the covariance form to the moment-based expression introduced by Kumar, Krackhardt, and Feld [Proc. Natl. Acad. Sci. 121, e2306412121 (2024)], which involves the (1)(-1)st, 11st, 22nd, and 33rd moments of the degree distribution. The two formulations are shown to be equivalent, as they should be: the moment-based representation expands the same structural dependence that the covariance form expresses in its most compact and interpretable form. The analysis thus unifies node-level and moment-level perspectives on the friendship paradox, offering both a pedagogically transparent derivation and a direct bridge to recent theoretical developments.
@@ I Introduction
Recent work by Kumar, Krackhardt, and Feld [2] introduced a moment-based decomposition that expresses the difference between these two averages in terms of degree moments and an edge-level correlation coefficient, the inversity $\rho$, which captures the correlation between origin degree and inverse destination degree across oriented edges. The present paper establishes a compact node-level covariance identity that expresses the difference between the alter-based and ego-based averages as a normalized covariance between degree and neighbor-averaged degree. The equivalence between the two formulations is shown explicitly, including a clear mapping between node-averaged quantities, edge-sampled expectations, and degree moments. This connection illuminates the roles of heterogeneity and mixing patterns [3] and provides a concise criterion for equality of the two variants.
@@ Acknowledgements.
The author sincerely appreciates Hang-Hyun Jo (조항현) for verifying the covariance formula. The author also thanks Eun Lee (이은), with whom he initially discussed the problem and explored a plan to extend the analysis, as well as Mi Jin Lee (이미진) for organizing an informal workshop dedicated to the discussion of “incomplete ideas” where the topic was further developed. This research was supported by the National Research Foundation of Korea (NRF) under the grant RS-2021-NR061247.